OpenStudy (anonymous):

in triangle PQR sin P=.5, sinR=.3, and r= 13. find the length of p.

terenzreignz (terenzreignz):

A law of Sines problem?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Use the law of sines itself :P \[\huge \frac{p}{\sin(P)}= \frac{r}{\sin(R)}\]

OpenStudy (anonymous):

okay so p/sin .5 = 13/sin .3

terenzreignz (terenzreignz):

No need for that, you have already been given sin(P) I mean \[\huge \sin(P) = 0.5\] etc

OpenStudy (anonymous):

okay so btw this is katherineregister im using my sisters account so i have p/0.5=13/.3

terenzreignz (terenzreignz):

Yes. And now, finish it.

OpenStudy (anonymous):

13(.5)/.3=p

terenzreignz (terenzreignz):

Yeah, whatever that is :)

OpenStudy (anonymous):

lol okay i have a alot of these to do and i just cant manage to get through them by myself:( i dont undcerstand why lol i know what i am doing as long as someone is guiding me haha okay so

OpenStudy (anonymous):

21.17

OpenStudy (anonymous):

thats the answer right?

terenzreignz (terenzreignz):

Approximately, yes.

OpenStudy (anonymous):

okay ill tag you in my next one